# Characterizing and Understanding HGNN Training on GPUs

### v3

@@ Abstract.
Owing to their remarkable representation capabilities for heterogeneous graph data, Heterogeneous Graph Neural Networks (HGNNs) have been widely adopted in many critical real-world domains such as recommendation systems and medical analysis. Prior to their practical application, identifying the optimal HGNN model parameters tailored to specific tasks through extensive training is a time-consuming and costly process. To enhance the efficiency of HGNN training, it is essential to characterize and analyze the execution semantics and patterns within the training process to identify performance bottlenecks. In this study, we conduct an in-depth quantification and analysis of two mainstream HGNN training scenarios, including single-GPU and multi-GPU distributed training. Based on the characterization results, we disclose the performance bottlenecks and their underlying causes in different HGNN training scenarios and provide optimization guidelines from both software and hardware perspectives.
@@ 1. Introduction
In recent years, Graph Neural Networks (GNNs) have demonstrated exceptional capabilities in representing and processing graph data in non-Euclidean spaces (Wu et al., 2020a; Xu et al., 2018; Zhou et al., 2020; Wu et al., 2020b). The early successes of GNNs are predominantly in the domain of homogeneous graphs (HomoGs), characterized by a single type of entity and adjacency relation (Kipf and Welling, 2017; Veličković et al., 2018; Hamilton et al., 2017). However, many real-world data in complex systems are more aptly represented as heterogeneous graphs (HetGs) which consist of multiple types of entities and relations embodied by various types of vertices and edges, respectively. In contrast to HomoGs, HetGs possess not only structural information inherent in graph data but also rich semantic information harbored in the relations (Shi et al., 2016; Wang et al., 2020a). Due to the powerful representation ability of HetGs, Heterogeneous Graph Neural Networks (HGNNs) have been developed and widely adopted in many critical fields including recommendation systems (Zhao and Yu, 2021; Altaf et al., 2019; xia Liang et al., 2022), cybersecurity (Hou et al., 2017; Ye et al., 2019; Fan et al., 2018), medical analysis (Luo et al., 2021; Kim et al., 2023), traffic prediction (Nazzal et al., 2024; Jin et al., 2021; Zhong et al., 2021) and many others.
@@
    Comparison between Backward and Forward Propagation: (1) Forward propagation is generally more time-consuming than backward propagation; (2) Although the kernels used during backward propagation are similar to those in forward propagation, there are significant differences in time distribution across kernel types and in the execution characteristics of the same kernels; (3) Backward propagation demands greater memory access and exhibits lower data locality than the forward pass. However, execution stages characterized by a high number of additions are anticipated to incur a lower computational workload in backward propagation.
@@
    Overall Analysis of End-to-end Training Process: (1) In comparison to pure inference, the training process necessitates greater memory footprint, with the majority memory allocation lies on the NA stage; (2) The vast majority of kernel stalls during the training process originate from memory dependency except for sgemm kernel; (3) The principal factor affecting the speedup of multi-GPU distributed training is the competition for various shared hardware resources.
@@ 2.1. Heterogeneous Graphs and Semantic Graphs
The semantic graphs are derived from the original heterogeneous graph based on specific relationships or metapaths. Each semantic graph contains only one type of relationship or metapath. As shown in Fig. 1, we build a semantic graph based on AP, PP, PS, and PSP, respectively. Semantic graphs separate multiple semantics in the original HetG, facilitating information extraction. Numerous HGNN models (Wang et al., 2019a; Fu et al., 2020; Yang et al., 2023) utilize metapaths to construct semantic graphs, while others (Schlichtkrull et al., 2018; Wang et al., 2020b) employ direct relations, i.e., edge types, for semantic graph construction.
@@ 2.3. HGNN Training
Fig. 2 depicts the training process of the HGNN models. The process of a single training epoch can generally be divided into four main steps as illustrated in Fig. 2(c). Firstly, during forward propagation, the embeddings of target vertices are computed according to the procedural steps as in model formula. Secondly, the loss computation (LC) stage transforms these embeddings into the vector space of classification categories, subsequently generating a probability distribution used to compute the loss function. Thirdly, backward propagation calculates the gradient of each model parameter relative to the loss function, employing the chain rule to determine the direction for parameter adjustments that yield the most rapid loss reduction. Finally, in the parameters update (PU) stage, model parameters are adjusted based on these gradients and a predefined learning rate.
Figure 2. Illustration of HGNN training: (a) SGB stage; (b) Mini-batch Sampling stage; (c) Training process on a single computing node; (d) Distributed training process.
@@ 3.1.1. Platforms
The configuration of our experimental platform is detailed in Table 1, featuring computing nodes equipped with four advanced A100 GPUs. The four GPUs are organized into two groups, with intra-group connections facilitated by NVLink buses, while inter-group connections are established through PCIe buses. We utilize Nsight Systems and Nsight Compute tools to capture detailed performance metrics data of execution on the GPUs. Regarding the software framework, we choose DGL (Wang et al., 2019b), which emerges as one of the most prominent GNN frameworks, typically outperforming PyG (Fey and Lenssen, 2019) in terms of runtime efficiency and energy consumption (Huang et al., 2022). All the experiments are conducted utilizing 32-bit floating-point data format.
@@ 3.1.3. Benchmark Datasets
We employ four widely-used HetG datasets as benchmark datasets: ACM, IMDB, DBLP, and OGBN-MAG (MAG), as detailed in Table 3. ACM, DBLP, and MAG are citation datasets, while IMDB represents a movie dataset. These datasets range in size from tens of thousands to tens of millions and billions of edges, providing a comprehensive representation of graph data across diverse real-world application scenarios.
@@ 3.2. Evaluation Methods
A fundamental and crucial step in profiling involves delineating the scope of the profiling process. We utilize the CUDA interface provided by PyTorch to initiate and terminate the cudaProfiler. Additionally, we employ NVTX tags to encapsulate the code regions designated for profiling, thereby differentiating between distinct execution phases.
Nsight Systems is utilized to provide a holistic analysis of system-level performance, offering insights into the entire application execution process, including both GPU and CPU activities, as well as their interactions with system resources. Meanwhile, Nsight Compute specializes in conducting detailed analysis specifically on the execution of CUDA kernels on the GPU. It offers granular performance metric data at the GPU instruction level, such as execution time, memory access patterns, and instruction efficiency for each CUDA kernel function.
Due to the GPU initialization processes involved in the initial training epochs, unless otherwise specified, the data presented in this paper are the geometric mean (GM) of results from the 5 epochs after excluding the first 3 epochs, which ensures the accuracy of the obtained results.
@@ 4.1. Full-batch Training
Full-batch training enables the model to access the entire graph for thorough parameter updates, leading to accurate and stable gradient estimates; however, it necessitates substantial memory resources. In this section, we present a detailed quantitative analysis of execution time, execution bounds, memory patterns, and instruction issue stalls during full-batch training. Additionally, we examine the influence of variations in metapath properties on execution performance. Due to an Out of Memory (OOM) issue encountered while conducting full-batch training on the MAG dataset with the A100 GPU, this section includes results solely for the three smaller datasets. Furthermore, as the SGB stage is executed only once throughout the entire training process, its associated analysis will be presented in Section 6.1.1.
@@ 4.1.1. Execution Time Analysis
Time Breakdown by Stage. Breaking down the execution time by stages allows us to understand the time proportion of each execution stage, thereby identifying the main execution stages for targeted optimization. Fig. 3 shows the profiling results of the time breakdown by stages. From the overall training perspective, 1 forward propagation is more time-consuming than backward propagation. This phenomenon occurs because, although the backward stage involves the reverse operations of forward propagation, it can directly reuse numerous intermediate results from the forward process. The experimental result in Fig. 3(a) indicates that forward propagation accounts for an average of 48.33% of the total execution time across different HGNN models and datasets, while backward propagation averages 42.37%. In terms of HGNN execution stages, 2 the NA stage dominates the most execution time both in forward and backward propagation, which is considered the predominant stage across all the HGNN stages. This is because each edge in every semantic graph requires one execution process during the NA stage, making the load during NA stage the heaviest. As shown in Fig. 3(b) and Fig. 3(c), the NA stage accounts for an average of 77.80% of the execution time in the forward propagation and 77.53% in the backward propagation.
Figure 3. Time breakdown of HGNN training by phase: (a) The whole training phase; (b) Forward; (c) Backward.
@@
The DM-Type kernels execute dense-dense matrix multiplication (DeMM) tasks, such as sgemm, and typically exhibit a regular execution pattern with a high compute-to-memory access ratio. The TB-Type kernels handle computational operations based on the irregular topologies of graphs, exemplified by operations like SpMMCsr and SDDMMCoo (sampled dense-dense matrix multiplication, SDDMM), often demonstrating an irregular execution pattern induced by the irregular neighbor connection patterns in graphs. The EW-Type kernels perform element-wise computational operations on sets of vectors or matrices, represented by elementwise_kernel (EleWise), matrix_scalar_kernel (MatScla), and reduce_kernel (Reduce), typically characterized by a low compute-to-memory access ratio. The DR-Type kernels specialize in data rearrangement tasks on matrices, such as CatArrayBatchedCopy (Concat) and DeviceRadixSortSingleTileKernel (Sort), involving a substantial amount of data movement operations.
@@
From an overall perspective, 3 the distribution of primary executing kernel types during the forward and backward propagation of the same stage exhibits significant similarity, whereas the time proportion of different kernel types varies. The resemblance in the distribution of kernel types stems from backward propagation being the inverse operation of forward propagation in model computations. The disparities in the distribution of kernel execution times originate from discrepancy in input data and computational loads during backward propagation. For example, as shown in Fig. 4(a), in the forward propagation, the FP stage is predominantly occupied by the DM-Type kernel (mainly sgemm), taking up an average of 96.48% of the total kernel execution time across various models and datasets. The EW-Type kernels (mainly EleWise) only occupy a small fraction which is less than 4%. However, during the backward propagation, the proportion of time spent on EW-Type kernels notably escalates, averaging 25.78% of the total kernel execution duration. The underlying cause of this phenomenon lies in the vertex-type-specific nature of the weight matrix employed during the FP stage which will be explained further in Section 6.1.2.
Note that the RGCN and RGAT models do not invoke any CUDA kernels during the backward propagation of SF stage as in Fig. 4(b). This is due to their adoption of a direct summation aggregation approach in the forward propagation of SF stage. Moreover, in a broader context, 4 addition operations during forward propagation necessitate no computational effort during in the backward pass. This is determined by the nature of the chain rule for gradient propagation, which can be explained by the following formula: $Loss=f(C)=f(A+B)\ \rightarrow\frac{\partial L}{\partial B}=\frac{\partial L}{\partial A}=\frac{\partial L}{\partial C}$.
@@ 4.1.2. Execution Bounds Analysis
In this section, we conduct a detailed analysis of the performance metrics of CUDA kernels predominantly invoked during each execution stage to identify the hardware resource bounds encountered at various stages.
Forward Propagation Analysis. The forward propagation process constitutes the primary execution phase of HGNN model training. Characterizing and analyzing the execution bounds during this phase is essential for understanding the requirements of various hardware resources throughout the model’s execution. Table 4 presents the performance metrics of the primary kernels at each execution stage during the training of the HAN model on the DBLP dataset. Based on the data in the table, we formulate the Roofline model shown in Fig. 5, which elucidates the execution bounds of each stage encountered during both forward and backward propagation.
@@
5
During the forward propagation, a distinct hybrid execution pattern emerges between stages, each characterized by unique execution bounds, resulting in varying demands on hardware resources. This is due to the distinct execution behaviors exhibited by different stages of the HGNN models. To be specific, for the FP stage, the kernels that occupy the majority of execution during forward propagtion is mainly sgemm, which primarily performs DeMM operations. As shown in Table 4, due to its regular memory access pattern, it achieves an L2 cache hit rate of 86.14%, while its high arithmetic density of 111.75 FLOP/Byte reflects its elevated computational demand. In Fig. 5, the sgemm kernel during the forward propagation is situated in the compute-bound region. The NA stage primarily involves graph-topology-based and element-wise operations, mainly invoking SpMMCsr and SDDMMCoo kernels as presented in Table 4. Taking the former as an example, its DRAM bandwidth utilization is 53.11%, but its arithmetic density is only 2.19 FLOP/Byte, indicating its high demand for memory resources. In Fig. 5, it lies within the memory-bound region. In the SF stage, the sgemm kernel is initially utilized for computing attention weights, subsequent to which EleWise and Reduce kernels are employed for aggregating features from diverse semantic graphs. According to the aforementioned analysis, this stage demonstrates an initial manifestation of compute-bound behavior, succeeded by memory-bound characteristic as depicted in Fig. 5.
Comparison of Backward and Forward. As a crucial component of the training process, backward propagation shares some similarities with but also exhibits significant differences from forward propagation. Contrasting them aids in a deeper understanding of the varying hardware resource demands during HGNN training. Overall, 6 the execution bounds in the backward and forward propagation exhibit similarity, yet the hybrid execution pattern in backward propagation is notably more intricate. This phenomenon arises because, while backward propagation reverses the process of forward propagation, the application of the chain rule in gradient propagation may introduce operations within the same stage that deviate from those encountered in forward propagation. As shown in Table 4, in FP-Backward compared to FP-Forward, the proportion of time spent on sgemm decreases from 98.94% to 45.40% , while the time occupied by EleWise kernels facing memory-bound notably increases. In both forward and backward propagation, the NA stage primarily exhibits memory-bound. However, what differs is that in the NA-Backward stage, the proportion of time taken by the SDDMMCoo kernel increases from 8.21% to 46.21%, which indicates the matrices corresponding to the operations performed in the NA stage during backward propagation are denser compared to those in the forward process. The execution bounds exhibited by the SF stage in both backward and forward propagation are fundamentally similar.

### v4

@@ Abstract.
Owing to their remarkable representation capabilities for heterogeneous graph data, Heterogeneous Graph Neural Networks (HGNNs) have been widely adopted in many critical real-world domains such as recommendation systems and medical analysis. Prior to their practical application, identifying the optimal HGNN model parameters tailored to specific tasks through extensive training is a time-consuming and costly process. To enhance the efficiency of HGNN training, it is essential to characterize and analyze the execution semantics and patterns within the training process to identify performance bottlenecks. In this study, we conduct a comprehensive quantification and in-depth analysis of two mainstream HGNN training scenarios, including single-GPU and multi-GPU distributed training. Based on the characterization results, we reveal the performance bottlenecks and their underlying causes in different HGNN training scenarios and propose optimization guidelines from both software and hardware perspectives.
@@ 1. Introduction
In recent years, Graph Neural Networks (GNNs) have demonstrated exceptional capabilities in representing and processing graph data in non-Euclidean spaces (Wu et al., 2020a; Xu et al., 2018; Zhou et al., 2020; Wu et al., 2020b). The early successes of GNNs are predominantly in the domain of homogeneous graphs (HomoGs), characterized by a single type of entity and adjacency relation (Kipf and Welling, 2017; Veličković et al., 2018; Hamilton et al., 2017). However, many real-world data in complex systems are more aptly represented as heterogeneous graphs (HetGs) which consist of multiple types of entities and relations embodied by various types of vertices and edges, respectively. In contrast to HomoGs, HetGs contain not only the structural information inherent in graph data but also the rich semantic information embedded in the relations (Shi et al., 2016; Wang et al., 2020a). Due to the powerful representation ability of HetGs, Heterogeneous Graph Neural Networks (HGNNs) have been developed and widely adopted in many critical fields including recommendation systems (Zhao and Yu, 2021; Altaf et al., 2019; xia Liang et al., 2022), cybersecurity (Hou et al., 2017; Ye et al., 2019; Fan et al., 2018), medical analysis (Luo et al., 2021; Kim et al., 2023), traffic prediction (Nazzal et al., 2024; Jin et al., 2021; Zhong et al., 2021) and many others.
@@
    Comparison between Backward and Forward Propagation: (1) Forward propagation is generally more time-consuming than backward propagation; (2) Although the kernels used during backward propagation are similar to those in forward propagation, there are significant differences in time distribution across kernel types and in the execution characteristics of the same kernels; (3) Backward propagation requires greater memory access and demonstrates lower data locality compared to the forward pass. However, stages of execution that involve a high number of additions are expected to have a lower computational workload during backward propagation.
@@
    Overall Analysis of End-to-end Training Process: (1) In comparison to pure inference, the training process necessitates greater memory footprint, with the majority memory allocation lies on the NA stage; (2) The vast majority of kernel stalls during the training process originate from memory dependency except for sgemm kernel; (3) The primary factor influencing the performance improvement of multi-GPU distributed training is the contention for shared hardware resources.
@@ 2.1. Heterogeneous Graphs and Semantic Graphs
Semantic graphs are derived from the original HetG by focusing on specific relations or metapaths. Each semantic graph is constructed to represent only a single type of relation or metapath. As illustrated in Fig. 1, semantic graphs are constructed based on the AP, PP, PS, and PSP, respectively. This process of segmentation allows for the separation of multiple semantic aspects within the original HetG, thereby facilitating more efficient information extraction. Numerous HGNN models (Wang et al., 2019a; Fu et al., 2020; Yang et al., 2023) utilize metapaths to construct semantic graphs, while others (Schlichtkrull et al., 2018; Wang et al., 2020b) employ direct relations, i.e., edge types, for semantic graph construction.
@@ 2.3. HGNN Training
Fig. 2 depicts the training process of the HGNN models. The process of a single training epoch can generally be divided into four main steps as illustrated in Fig. 2(c). Firstly, during forward propagation, the embeddings of target vertices are computed according to the procedural steps as in the model formula. Secondly, the loss computation (LC) stage transforms these embeddings into the vector space of classification categories, subsequently generating a probability distribution used to compute the loss function. Thirdly, backward propagation calculates the gradient of each model parameter relative to the loss function, employing the chain rule to determine the direction for parameter adjustments that yield the most rapid loss reduction. Finally, in the parameters update (PU) stage, model parameters are adjusted based on these gradients and a predefined learning rate.
Figure 2. Illustration of HGNN training: (a) SGB stage; (b) Mini-batch sampling process; (c) Training process on a single computing node; (d) Distributed training process.
@@ 3.1.1. Platforms
The configuration of our experimental platform is detailed in Table 1, featuring computing nodes equipped with four advanced A100 GPUs. The four GPUs are organized into two groups, with intra-group connections facilitated by NVLink buses, while inter-group connections are established through PCIe buses. We utilize Nsight Systems and Nsight Compute tools to capture detailed performance metrics data of execution. Regarding the software framework, we choose DGL (Wang et al., 2019b), which emerges as one of the most prominent GNN frameworks, typically outperforming PyG (Fey and Lenssen, 2019) in terms of runtime efficiency and energy consumption (Huang et al., 2022). All the experiments are conducted utilizing 32-bit floating-point data format.
@@ 3.1.3. Benchmark Datasets
We employ four widely-used HetG datasets as benchmark datasets: ACM, IMDB, DBLP, and OGBN-MAG (MAG), as detailed in Table 3. ACM, DBLP, and MAG are citation datasets focused on academic publications, capturing intricate relationships between papers, authors, and institutions, making them well-suited for tasks such as citation analysis and academic recommendation systems. IMDB is a movie dataset containing detailed information about films, actors, and directors, commonly employed in movie recommendation systems and social network analysis. Ranging in size from tens of thousands to billions of edges, they offer comprehensive graph representations across a wide range of real-world applications.
@@ 3.2. Evaluation Methods
A fundamental and critical step in profiling involves defining the scope of the profiling process. We utilize the CUDA interface provided by PyTorch to initiate and terminate the cudaProfiler. Furthermore, NVTX tags are employed to demarcate the code regions designated for profiling, thereby distinguishing between different execution phases.
Nsight Systems is employed to deliver a comprehensive analysis of system-level performance, offering insights into the complete application execution process, encompassing both GPU and CPU activities, as well as their interactions with system resources. In contrast, Nsight Compute is dedicated to providing an in-depth analysis of CUDA kernel executions on the GPU. It supplies detailed performance metrics at the GPU instruction level, including execution time, memory access patterns, and instruction efficiency for each CUDA kernel function.
Due to the GPU initialization processes involved in the initial training epochs, unless otherwise specified, the data presented in this paper are the geometric mean (GM) of results from the 5 epochs after excluding the first 3 epochs, which ensures the accuracy and reliability of the obtained results.
@@ 4.1. Full-batch Training
Full-batch training enables the model to access the entire graph for thorough parameter updates, leading to accurate and stable gradient estimates; however, it necessitates substantial memory resources. In this section, we provide a detailed quantitative analysis of execution time, execution bounds, memory access patterns, and instruction issue stalls during full-batch training. Furthermore, we investigate the impact of variations in metapath properties on execution performance. Due to an Out of Memory (OOM) issue encountered while conducting full-batch training on the MAG dataset with the A100 GPU, this section includes results solely for the three smaller datasets. Furthermore, as the SGB stage is executed only once throughout the entire training process, its associated analysis will be presented in Section 6.1.1.
@@ 4.1.1. Execution Time Analysis
Figure 3. Time breakdown of HGNN training by phase: (a) The whole training process; (b) Forward; (c) Backward.
Time Breakdown by Stage. Breaking down the execution time by stages allows us to understand the time proportion of each execution stage, thereby identifying the main execution stages for targeted optimization. Fig. 3 shows the profiling results of the time breakdown by stages. From the overall training perspective, 1 forward propagation is more time-consuming than backward propagation. This phenomenon occurs because, although the backward stage involves the reverse operations of forward propagation, it can directly reuse numerous intermediate results from the forward process. The experimental result in Fig. 3(a) indicates that forward propagation accounts for an average of 48.33% of the total execution time across different HGNN models and datasets, while backward propagation averages 42.37%. In terms of HGNN execution stages, 2 the NA stage dominates the most execution time both in forward and backward propagation, which is considered the predominant stage across all the HGNN stages. This is due to the fact that each edge in every semantic graph necessitates a distinct execution process during the NA stage, making the load during NA stage the heaviest. As shown in Fig. 3(b) and Fig. 3(c), the NA stage accounts for an average of 77.80% of the execution time in the forward propagation and 77.53% in the backward propagation.
@@
The DM-Type kernels perform DeMM tasks, such as sgemm, and typically exhibit a regular execution pattern with a high compute-to-memory-access ratio. The TB-Type kernels manage computational operations based on the irregular topologies of graphs, exemplified by operations like SpMMCsr and SDDMMCoo (sampled dense-dense matrix multiplication, SDDMM), and often demonstrate an irregular execution pattern due to the uneven neighbor connection patterns in graphs. The EW-Type kernels execute element-wise computational operations on sets of vectors or matrices, represented by elementwise_kernel (EleWise), matrix_scalar_kernel (MatScla), and reduce_kernel (Reduce), typically characterized by a low compute-to-memory-access ratio. The DR-Type kernels are specialized in data rearrangement tasks on matrices, such as CatArrayBatchedCopy (Concat) and DeviceRadixSortSingleTileKernel (Sort), and involve a significant amount of data movement operations.
@@
From an overall perspective, 3 the distribution of primary executing kernel types during the forward and backward propagation of the same stage exhibits significant similarity, whereas the proportion of time consumed by these kernel types varies between the forward and backward propagation. The resemblance in the distribution of kernel types stems from backward propagation being the inverse operation of forward propagation in model computations. The disparities in the distribution of kernel execution times originate from discrepancy in input data and computational loads during backward propagation. For example, as shown in Fig. 4(a), in the forward propagation, the FP stage is predominantly occupied by the DM-Type kernel (mainly sgemm), taking up an average of 96.48% of the total kernel execution time across various models and datasets. The EW-Type kernels (mainly EleWise) only occupy a small fraction which is less than 4%. However, during the backward propagation, the proportion of time spent on EW-Type kernels notably escalates, averaging 25.78% of the total kernel execution duration. The underlying cause of this phenomenon lies in the vertex-type-specific nature of the weight matrix employed during the FP stage which will be further explained in Section 6.1.2.
Note that the RGCN and RGAT models do not invoke any CUDA kernels during the backward propagation of SF stage as in Fig. 4(b). This is due to their adoption of a direct summation aggregation approach in the forward propagation of SF stage. Moreover, in a broader context, 4 addition operations during forward propagation necessitate no computational effort during the backward pass, and this phenomenon is observed in every execution stage. For instance, during the projection process in the FP stage, linear transformations typically involve the addition of a bias term, applied to the resulting matrix through element-wise addition. Additionally, in the NA stage, a significant amount of element-wise vector addition is used to aggregate the feature vectors of neighboring vertices, with a similar operation occurring in the SF stage. Consider matrices, vectors of the same dimensions, or scalars denoted as $A$, $B$, and $C$, with the relation $A=B+C$. If the forward propagation loss is defined as $Loss=f(A)=f(B+C)$, given the gradient of $A$ with respect to $Loss$: $\frac{\partial L}{\partial A}$, according to the chain rule of gradient propagation, since the partial derivatives of $A$ with respect to $B$ and $C$ are both 1, the gradients of $B$ and $C$ can be directly obtained from $\frac{\partial L}{\partial A}$. Specifically, the gradients of $B$ and $C$ are equal to the gradient of $A$, without any additional computation. This rule can be formally represented by the equation: $Loss=f(A)=f(B+C)\ \rightarrow\frac{\partial L}{\partial B}=\frac{\partial L}{\partial C}=\frac{\partial L}{\partial A}$.
@@ 4.1.2. Execution Bounds Analysis
In this section, we perform a comprehensive analysis of the performance metrics of CUDA kernels predominantly invoked at each execution stage, with the aim of identifying the hardware resource bounds encountered throughout the various stages.
Forward Propagation Analysis. The forward propagation process constitutes the primary execution phase of HGNN model training. Characterizing and analyzing the execution bounds during this phase is essential for understanding the requirements of various hardware resources throughout the model’s execution. Table 4 presents the performance metrics of the primary kernels at each execution stage during the training of the HAN model on the DBLP dataset. Based on the data in the table, we formulate the Roofline model (Williams et al., 2009) shown in Fig. 5, which elucidates the execution bounds of each stage encountered during both forward and backward propagation.
5
During the forward propagation, a distinct hybrid execution pattern emerges between stages, each characterized by unique execution bounds, resulting in varying demands on hardware resources. This is due to the distinct execution behaviors exhibited by different stages of the HGNN models. To be specific, for the FP stage, the kernels that occupy the majority of execution during forward propagtion is mainly sgemm, which primarily performs DeMM operations. As shown in Table 4, due to its regular memory access pattern, it attains an L2 cache hit rate of 86.14%, while its high arithmetic density of 111.75 FLOP/Byte indicates a substantial computational demand. In Fig. 5, the sgemm kernel during the forward propagation is situated in the compute-bound region. The NA stage primarily involves graph-topology-based and element-wise operations, mainly invoking SpMMCsr and SDDMMCoo kernels as presented in Table 4. Taking the former as an example, its DRAM bandwidth utilization is 53.11%, but its arithmetic density is only 2.19 FLOP/Byte, indicating its high demand for memory resources. In Fig. 5, it lies within the memory-bound region. In the SF stage, the sgemm kernel is initially utilized for computing attention weights, subsequent to which EleWise and Reduce kernels are employed for aggregating features from diverse semantic graphs. According to the aforementioned analysis, this stage demonstrates an initial manifestation of compute-bound behavior, succeeded by memory-bound characteristic as depicted in Fig. 5.
@@
Comparison of Backward and Forward Propagation. As a crucial component of the training process, backward propagation shares some similarities with but also exhibits significant differences from forward propagation. Contrasting them aids in a deeper understanding of the varying hardware resource demands during HGNN training. Overall, 6 the execution bounds in the backward and forward propagation exhibit similarity, yet the hybrid execution pattern in backward propagation is notably more intricate. This phenomenon arises because, while backward propagation reverses the process of forward propagation, the application of the chain rule in gradient propagation may introduce operations within the same stage that deviate from those encountered in forward propagation. As shown in Table 4, in FP-Backward compared to FP-Forward stage, the proportion of time spent on sgemm decreases from 98.94% to 45.40%, while the time occupied by EleWise kernels facing memory-bound notably increases. In both forward and backward propagation, the NA stage primarily exhibits memory-bound. However, what differs is that in the NA-Backward stage, the proportion of time taken by the SDDMMCoo kernel increases from 8.21% to 46.21%, which indicates the matrices involved in matrix multiplication during the NA stage of backward propagation are denser compared to those in forward propagation. This is due to the fact that, during backward propagation of NA stage, in addition to gradient propagation through edges, it is necessary to compute gradients of the vertex features first, which involves operating on dense matrices. The execution bounds exhibited by the SF stage in both backward and forward propagation are fundamentally similar.